OpenStudy (anonymous):

1,440 is 120 percent of

OpenStudy (anonymous):

Let 120 percent of what is x: So: \[120 \; \% of \; x = 1440\]

OpenStudy (anonymous):

\[\frac{120}{100} \times x = 1440\] \[x = 1440 \times \frac{100}{120} = ??\]

OpenStudy (kaederfds):

100/120 x 1440 = 1400

OpenStudy (kaederfds):

is that true?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

No that is not true..

OpenStudy (anonymous):

1,140.63 1,140.62 1,104.63 1,014.62

OpenStudy (kaederfds):

1200

OpenStudy (anonymous):

Solve that..

OpenStudy (anonymous):

\[1440 \times \frac{100}{120} = ??\]

OpenStudy (anonymous):

17280000 1440x100x120????

OpenStudy (anonymous):

\(1440 \times 100 = ??\)

OpenStudy (anonymous):

144000

OpenStudy (anonymous):

Now : \[\frac{144000}{120} = ??\]

OpenStudy (anonymous):

Are those 4 you have written your answer choices ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then check for question properly they can't if this is your question..